2 goods (perfect complements) production function: f(L) = V30L . Fixed cost = 3. Find the supply function.
Answer #1

q= f(L) = (30L)^0.5

Squaring both sides

q^2 = 30L

L= 30/q^2

Variable cost= VC= wL = 30w/q^2

Marginal cost is the supply curve: MC= P

MC= d(VC)/dq= -60w/q^3

P= -60w/q^3

q= (-60w/P)^1/3

This is the supply function.